Use implicit differentiation to find dy/dx:
e^y lnx+ xy3= y^2

Answer by venugopalramana(3286)

e^y lnx+ xy3= y^2...DIFFERENTIATING WRT X
(E^Y)(DY/DX)(LN(X))+(1/X)(E^Y)+Y^3+X(3Y^2)(DY/DX)=2Y(DY/DX)
(DY/DX){(E^Y)(LN(X))+X(3Y^2)-2Y}=-{(1/X)(E^Y)+Y^3}
DY/DX=-{(1/X)(E^Y)+Y^3}/{(E^Y)(LN(X))+X(3Y^2)-2Y}